# Blind Estimation of Sub-band Acoustic Parameters from Ambisonics Recordings using Spectro-Spatial Covariance Features

### v1

@@ II-A Spectro-Spatial Covariance Vector (SSCV)
To effectively capture the spatial, spectral and temporal properties of sound from FOA signals, we developed the SSCV feature. The feature is designed to encode both the time-frequency power distribution, representing spectral information, and the inter-channel correlation, including the phase differences between channels. This dual capability allows SSCV to provide a comprehensive representation of spatial audio characteristics, which is crucial for acoustic parameter estimation tasks. The extraction process, depicted in Figure 2, consists of three key stages: banded covariance computation, smoothing, and vectorization.
@@ III-C Training and Evaluation
All models involved in this study were trained for 100 epochs using the Adam optimizer [26] with an initial learning rate of 0.0005, and dropout of $20\%$ on the output of the convolutional block. We halved the learning rate if the validation loss did not decrease for 5 consecutive epochs. The batch size was set to 128. For each parameter, optimization was performed using the mean square error (MSE) for all sub-bands as the loss function. For the T60 estimation in our proposed model, the logarithm of both the reference and estimated T60 values were used when calculating the MSE, as psychoacoustic studies indicate that humans perceive T60 on a logarithmic scale [27]. The model with the lowest validation loss was selected for evaluation.

### v2

@@ II-A Spectro-Spatial Covariance Vector (SSCV)
The effective representation of multi-channel speech recordings has been extensively studied in tasks like target speaker extraction [22] and room volume estimation [17]. Prior research highlights the importance of incorporating time-frequency power distribution and inter-channel differences for robust performance. To address this need in FOA signals, we developed the SSCV feature, designed to encode both time-frequency power distribution and inter-channel correlation, including phase differences. This dual capability provides a comprehensive representation of spatial audio, crucial for acoustic parameter estimation. The extraction process, depicted in Figure 2, involves three stages: banded covariance computation, smoothing, and vectorization.
@@ III-C Training and Evaluation
All models in this study were initially set to train for 100 epochs using the Adam optimizer [27] with an initial learning rate of 0.0005 and a $20\%$ dropout rate applied to the output of the convolutional block. The learning rate was halved if the validation loss did not improve for five consecutive epochs, and early stopping was employed if the validation loss did not decrease for 10 epochs. The batch size was set to 128. For each parameter, optimization was performed using the mean square error (MSE) for all sub-bands as the loss function. For the T60 estimation in our proposed model, the logarithm of both the reference and estimated T60 values were used when calculating the MSE, as psychoacoustic studies indicate that humans perceive T60 on a logarithmic scale [28]. The model with the lowest validation loss was selected for evaluation.